OpenStudy (zappy620):

How to find the derivative of xsinx where x=pi/3

OpenStudy (zappy620):

The pi/3 has got me confused :(

OpenStudy (zappy620):

\[\frac{ d }{ dx }[ x \sin (x)] ,x= \frac{ \pi }{ 3 }\]

OpenStudy (triciaal):